Grade 12th passPhysical Chemistry

1 A sample of FeSO4 and FeC2O4 dissolved in H2SO4. the complete oxidation of sample required8./3 eq of KMnO4 After oxdation the reaction mixture Was reduced by Zn On again 0Xidation byKMmO4Required 5/3eq the mole ratio of FeSO4 and FeC2O4

mehrose
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago
equivalent of kmno4= equivalent of fe+2 + equivalent of c2o4-2 + equivalent of fe+2

8/3=2x+equivalent of fe+2

equivalent of kmno4 = equivalent of fe+3 formed

5/3= x+ equivalent of fe+3

let equivalent of feso4=y

then equivalent of fe+3 formed = y

so

8/3 = 2x+y

5/3=x+y

x=1

and y=2/3

equivalent of feso4 = 2/3 and equivalent of fec2o4=1

moles*f=2/3

moles=1/3

moles*f=1

moles=1/2

moles ratio = 2/3